# Affleck-Dine leptogenesis scenario for resonant production of sterile neutrino dark matter

### v2

@@ 2.2 Numerical setup
$\frac{{\rm{d}}}{{\rm{d}}t}L_{\nu_{e}}(T)=\frac{45}{4\pi^{4}g_{*,s}(T_{i})}\int{\rm{d}\epsilon}\;\epsilon^{2}\Gamma_{\nu_{e}}(p,T)\left[\theta_{M}^{2}(\epsilon,T)f_{\nu_{e}}(\epsilon,T)-\bar{\theta}_{M}^{2}(\epsilon,T)f_{\bar{\nu}_{e}}(\epsilon,T)\right]\ .$ (2.19)
@@ 3.1 Affleck-Dine leptogenesis
Now, we comment on the flat direction considered in our scenario. As a concrete flat direction, we consider $\mu^{c}L_{e}L_{\tau}$ or $\tau^{c}L_{e}L_{\mu}$ direction, where $a^{c}$ is the charge conjugation of an SU(2)${}_{\rm{W}}$ singlet charged lepton with flavor $a=e,\mu,\tau$, and $L_{a}$ is an SU(2)${}_{\rm{W}}$ doublet lepton [38]. As a result, the generated lepton asymmetry satisfies $L_{\nu_{e}}=-L_{\nu_{\mu}}=L_{\nu_{\tau}}$ or $L_{\nu_{e}}=L_{\nu_{\mu}}=-L_{\nu_{\tau}}$, respectively. Then, the first term in the right-hand side in Eq. (2.16) is equal to $2\sqrt{2}G_{\rm{F}}L_{\nu_{e}}s_{\rm{tot}}$, which can induce the resonant production of sterile neutrinos. [^2] On the other hand, $L=0$ for $e^{c}L_{\mu}L_{\tau}$ flat direction.
@@ 3.2 Delayed-type L-ball scenario
where $\beta_{\rm{G}}\simeq 6\times 10^{-5}$ [45] in the case with $|\delta|\lesssim 10^{-1}$, which we consider in this paper, and we adopt $\zeta\simeq 2.4$ [46]. The L-balls decay into neutrinos with the decay rate $\Gamma_{Q}$ given by [47, 48]
@@
Finally, we discuss the condition for L-balls to dominate the universe before the L-ball decay. To this end, we define a ratio $f_{\rm{Q}}$ of the energy density of L-balls (or the AD field) $\rho_{\rm{Q}}$ to that of the inflaton and inflaton decay products $\rho_{\rm{R}}$. Considering that $f_{\rm{Q}}$ is almost constant during $t_{\rm{osc}}\lesssim t\lesssim t_{\rm{R}}$, where $t_{\rm{R}}$ is the cosmic time at the completion of reheating, and scales as $\propto a$ after reheating, we obtain
$f_{\rm{Q}}(T_{\rm{decay}})\simeq\frac{g_{*}(T_{\rm{decay}})T_{\rm{decay}}^{4}}{g_{*}(T_{\rm{relic,dec}})T_{\rm{relic,dec}}^{4}}\simeq\frac{m_{3/2}^{2}\varphi_{\rm{osc}}^{2}}{3M_{\rm{Pl}}^{2}H_{\rm{osc}}^{2}}\frac{a_{\rm{decay}}}{a_{\rm{R}}}\ ,$ (3.15)
where $T_{\rm{relic,dec}}$ is the temperature of relic plasma from the inflaton right before the L-ball decay. From the entropy conservation, we obtain
$\frac{T_{\rm{relic,dec}}}{T_{\rm{R}}}\simeq\left(\frac{g_{*}(T_{\rm{R}})}{g_{*}(T_{\rm{relic,dec}})}\right)^{1/3}\frac{a_{\rm{R}}}{a_{\rm{decay}}}\ .$ (3.16)
Combining Eqs. (3.15) and (3.16), we obtain $f_{\rm{Q}}(T_{\rm{decay}})$ as
$f_{\rm{Q}}(T_{\rm{decay}})\simeq\frac{g_{*}(T_{\rm{R}})^{4/9}}{g_{*}(T_{\rm{decay}})^{1/3}g_{*}(T_{\rm{relic,dec}})^{1/9}}\left(\frac{T_{\rm{R}}}{T_{\rm{decay}}}\right)^{4/3}\left(\frac{\varphi_{\rm{osc}}^{2}}{3M_{\rm{Pl}}^{2}}\right)^{4/3}\ .$ (3.17)
@@
$M_{\rm{F}}\lesssim$ $~{}2.8\times 10^{6}\,{\rm{GeV}}\left(\frac{g_{*}(T_{\rm{decay}})}{80}\right)^{-1/8}\left(\frac{g_{*}(T_{\rm{low}})}{60}\right)^{1/10}\left(\frac{|\delta|}{10^{-3}}\right)^{1/10}$
@@
Next, we consider the baryon asymmetry produced from the evaporation process of L-balls until the sphaleron process becomes out of equilibrium [49]. The total evaporated baryon charge can be written as [43] [^3] Here, we assume that the reheating temperature $T_{\rm{R}}$ is less than $T_{\rm{eq}}$.
$\frac{\Delta Q_{b}}{Q_{G}}$ $\simeq-\frac{128\sqrt{2}}{207}\pi^{2}\zeta^{-1}\frac{M_{\rm{Pl}}T_{\rm{R}}^{2}}{M_{s}^{2}M_{F}}|Q_{G}|^{-3/4}\left(1+\log\frac{T_{\rm{eq}}}{T_{\rm{R}}}+\frac{27}{16\pi}\sqrt{\frac{10}{g_{*}}}\right)\ ,$ (3.20)
where $M_{s}$ is the mass of gauginos involved in the evaporation process. Here, $T_{\rm{eq}}\sim(M_{s}^{2}M_{\rm{F}})^{1/3}Q_{G}^{-1/12}$ is the temperature at which the evaporation rate of L-balls [49] and diffusion rate of leptons evaporated from L-balls [50] becomes equal. Then, the total baryon asymmetry induced from the L-ball evaporation process can be evaluated as
$\eta_{b}$ $\simeq\eta_{L}\left(\frac{\Delta Q_{b}}{Q_{\rm{G}}}\right)\sim-\frac{128\sqrt{2}}{207}\pi^{2}\zeta^{-1}\eta_{L}\frac{M_{\rm{Pl}}T_{\rm{R}}^{2}}{M_{s}^{2}M_{\rm{F}}}|Q_{\rm{G}}|^{-\frac{3}{4}}$
$\simeq-2.2\times 10^{-11}\left(\frac{g_{*}}{80}\right)^{-\frac{1}{4}}\left(\frac{\delta}{10^{-3}}\right)\left(\frac{M_{s}}{10^{4}\,{\rm{GeV}}}\right)^{-2}\left(\frac{m_{3/2}}{1\,{\rm{GeV}}}\right)^{\frac{9}{2}}\left(\frac{M_{\rm{F}}}{10^{6}\,{\rm{GeV}}}\right)^{-6}\left(\frac{T_{\rm{R}}}{10\,{\rm{GeV}}}\right)^{2},$ (3.21)
where $\eta_{L}$ is the lepton-to-entropy ratio and we take $\eta_{L}=L_{\nu_{e}}^{\mathrm{init}}$. Here, we neglected the factor $1+\log\frac{T_{\rm{eq}}}{T_{\rm{R}}}+\frac{27}{16\pi}\sqrt{\frac{10}{g_{*}}}\,(\sim\mathcal{O}(1))$ in the right-hand side in Eq. (3.20). We require that this resulting baryon asymmetry does not spoil the BBN, that is, $|\eta_{b}|<|\eta_{b}^{\rm{obs}}|\sim 10^{-10}$, which leads to the upper bound of $T_{\rm{R}}$ in terms of $M_{s},m_{3/2}$, and $M_{\rm{F}}$. On the other hand, the L-ball domination ($f_{\mathrm{Q}}\gg 1$) requires $T_{\mathrm{R}}\gtrsim T_{\mathrm{decay}}$. Thus, using Eq. (3.14) and the upper bound on $T_{\mathrm{R}}$, we obtain
$M_{\rm{F}}\gtrsim 6.4\times 10^{5}\,{\rm{GeV}}\left(\frac{g_{*}(T_{\rm{decay}})}{80}\right)^{-3/40}\left(\frac{|\delta|}{10^{-3}}\right)^{1/10}\left(\frac{m_{3/2}}{1\,{\rm{GeV}}}\right)^{19/20}\left(\frac{M_{s}}{10^{4}\,{\rm{GeV}}}\right)^{-1/5}.$ (3.22)
@@
Figure 6: Constraints on $m_{3/2}$ and $M_{\mathrm{F}}$ in the delayed-type L-ball scenario. We set $M_{s}=10\,{\rm{TeV}}$ in the upper panels and $M_{s}=10^{2}\,{\rm{TeV}}$ in the lower panels. We set $\delta=10^{-2}$ and $\delta=10^{-3}$ in the left and right panels, respectively. The green, violet, and brown lines correspond to $|L_{\nu_{e}}^{\mathrm{init}}|=10^{-4}$, $10^{-3}$, and $5\times 10^{-3}$, respectively. The red-shaded region is excluded because baryon asymmetry is overproduced, $|\eta_{b}|>10^{-10}$, due to the evaporation process of L-balls when we require $T_{\rm{R}}>T_{\rm{decay}}$ (see Eq. (3.22)). The cyan region is excluded because $T_{\rm{decay}}$ is lower than $T_{\rm{low}}$ (see Eq. (3.19)). For the red regions, we used $g_{*}=106.75$ in Eq. (3.20) as a typical value above the electroweak scale.
@@ 4.1 Free-streaming length of sterile neutrinos
where $m_{\rm{eff}}=94\,{\rm{eV}}\times\Omega_{\rm{DM}}h^{2}$ is the total mass of the active neutrinos for their energy density to be the same as dark matter. Using the relation between $m_{x}$ and $\lambda_{\mathrm{FS}}$, we translate the lower bound, $m_{x}>4.09$ keV ($95\%$ CL) [52], into the upper bound of $\lambda_{\rm{FS}}$, $\lambda_{\rm{FS}}<\lambda_{\rm{FS}}^{\rm{upper}}=0.76$ Mpc.
@@
where we used $p=\epsilon_{\rm{phys}}(T_{*})T_{*}a_{*}/a$. Now, we calculate the free-streaming length of the sterile neutrinos in the SF mechanism using the distribution function obtained in Sec. 2. We show the free-streaming length of the sterile neutrinos as a function of $m_{s}$ for a given $L_{\nu_{e}}^{\rm{init}}$ in Fig. 7. By equating $\lambda_{\rm{FS}}(m_{s},L_{\nu_{e}}^{\rm{init}})$ and $\lambda_{\rm{FS}}^{\rm{upper}}=0.76$ Mpc, we obtain the Lyman-$\alpha$ constraint on $m_{s}$ for a given $L_{\nu_{e}}^{\rm{init}}$. We show the lower bounds of $m_{s}$ for several values of $L_{\nu_{e}}^{\rm{init}}$ in Table 1.
@@ 4.2 Gravitational wave enhancement at L-ball decay
A 0.5 $1.3\times 10^{6}$ $1.0\times 10^{2}$ $1.7\times 10^{4}$ $8.7\times 10^{-11}$ $4.0\times 10^{-4}$
B 0.4 $1.0\times 10^{6}$ $1.0\times 10^{3}$ $2.0\times 10^{5}$ $8.7\times 10^{-11}$ $4.8\times 10^{-4}$
C 1.0 $2.9\times 10^{6}$ $1.5\times 10^{4}$ $6.6\times 10^{5}$ $8.7\times 10^{-11}$ $2.2\times 10^{-4}$
Table 2: Benchmark parameters to evalute the gravitational wave spectra.

### v3

@@ 2.2 Numerical setup
$\frac{{\rm{d}}}{{\rm{d}}t}L_{\nu_{e}}(T)=-\frac{45}{4\pi^{4}g_{*,s}(T_{i})}\int{\rm{d}\epsilon}\;\epsilon^{2}\Gamma_{\nu_{e}}(p,T)\left[\theta_{M}^{2}(\epsilon,T)f_{\nu_{e}}(\epsilon,T)-\bar{\theta}_{M}^{2}(\epsilon,T)f_{\bar{\nu}_{e}}(\epsilon,T)\right]\ .$ (2.19)
@@ 2.4 Numerical results
At the end of this section, we mention other proposed mechanisms for sterile neutrino dark matter production [35, 36]. In the model discussed in Ref. [35], the sterile neutrinos are produced via the decay of gauge singlet Higgs bosons at $T\sim 100$ GeV. In the model discussed in Ref. [36], sterile neutrinos have a gauged $U(1)_{B-L}$ charge, and they are mainly produced from standard model particles in thermal plasma via the $U(1)_{B-L}$ gauge boson exchange. In these models, sterile neutrinos are produced at much higher cosmic temperatures than in the resonant production model, and the averaged momentum of dark matter is more redshifted due to the time evolution of $g_{*}$. Thus, these scenarios can account for sterile neutrino dark matter with masses of a few keV evading the constraint from the observations of the Lyman-$\alpha$ forests.
@@ 3.1 Affleck-Dine leptogenesis
Now, we comment on the flat direction considered in our scenario. As a concrete flat direction, we consider $\mu^{c}L_{e}L_{\tau}$ or $\tau^{c}L_{e}L_{\mu}$ direction, where $a^{c}$ is the charge conjugation of an SU(2)W singlet charged lepton with flavor $a=e,\mu,\tau$, and $L_{a}$ is an SU(2)W doublet lepton [40]. As a result, the generated lepton asymmetry satisfies $L_{\nu_{e}}=-L_{\nu_{\mu}}=L_{\nu_{\tau}}$ or $L_{\nu_{e}}=L_{\nu_{\mu}}=-L_{\nu_{\tau}}$, respectively. Then, the first term in the right-hand side in Eq. (2.16) is equal to $2\sqrt{2}G_{\rm{F}}L_{\nu_{e}}s_{\rm{tot}}$, which can induce the resonant production of sterile neutrinos. [^2] On the other hand, $L=0$ for $e^{c}L_{\mu}L_{\tau}$ flat direction.
@@ 3.2 Delayed-type L-ball scenario
where $\beta_{\rm{G}}\simeq 6\times 10^{-5}$ [47] in the case with $|\delta|\lesssim 10^{-1}$, which we consider in this paper, and $\zeta$ is a numerical factor $\simeq 2^{1/4}(c_{0}/\pi)^{1/2}$ where $c_{0}\equiv 4.8\ln(m_{\phi}/\sqrt{2}\omega_{Q})+7.4$ [48]. In the following, we adopt $\zeta=3.4$, which is obtained by substituting $m_{3/2}=1$ GeV, $M_{\rm{F}}=10^{6}$ GeV and $m_{\phi}=10^{4}$ GeV. The L-balls decay into neutrinos with the decay rate $\Gamma_{Q}$ given by [49, 50]
@@
Finally, we discuss the condition for L-balls to dominate the universe before the L-ball decay. To this end, we define a ratio $f_{\rm{Q}}$ of the energy density of L-balls (or the AD field) $\rho_{\rm{Q}}$ to that of the inflaton and inflaton decay products $\rho_{\rm{R}}$. Considering that $f_{\rm{Q}}$ is almost constant during $t_{\rm{osc}}\lesssim t\lesssim t_{\rm{R}}$, where $t_{\rm{R}}$ is the cosmic time at the completion of reheating, and scales as $\propto T^{-1}$ after reheating, we obtain
$f_{\rm{Q}}(T_{\rm{decay}})\simeq\frac{g_{*}(T_{\rm{decay}})T_{\rm{decay}}^{4}}{g_{*}(T_{\rm{relic,dec}})T_{\rm{relic,dec}}^{4}}\simeq\frac{m_{3/2}^{2}\varphi_{\rm{osc}}^{2}}{3M_{\rm{Pl}}^{2}H_{\rm{osc}}^{2}}\frac{T_{\mathrm{R}}}{T_{\mathrm{relic,dec}}}\ ,$ (3.15)
where $T_{\rm{relic,dec}}$ is the temperature of relic plasma from the inflaton right before the L-ball decay, and $T_{\rm{R}}$ is the reheating temperature. Then, we obtain
$f_{\rm{Q}}(T_{\rm{decay}})\simeq\left(\frac{g_{*}(T_{\rm{relic,dec}})}{g_{*}(T_{\rm{decay}})}\right)^{1/3}\left(\frac{T_{\rm{R}}}{T_{\rm{decay}}}\right)^{4/3}\left(\frac{\varphi_{\rm{osc}}^{2}}{3M_{\rm{Pl}}^{2}}\right)^{4/3}\ .$ (3.16)
@@
$M_{\rm{F}}\lesssim$ $~{}3.1\times 10^{6}\,{\rm{GeV}}\left(\frac{g_{*}(T_{\rm{decay}})}{80}\right)^{-1/8}\left(\frac{g_{*}(T_{\rm{low}})}{60}\right)^{1/10}\left(\frac{|\delta|}{10^{-3}}\right)^{1/10}$
@@
Next, we consider the baryon asymmetry produced from the evaporation process of L-balls until the sphaleron process becomes out of equilibrium [51]. The total evaporated baryon charge can be written as [45] [^3] Here, we assume $T_{\rm{ew}}<T_{\rm{eq}}$.
$\frac{\Delta Q_{b}}{Q_{\rm G}^{\mathrm{init}}}\simeq\begin{cases}-\frac{16\pi}{23\sqrt{5}}\frac{\sqrt{g_{*}(T_{\rm{R}})}}{g_{*}(T_{\rm{ew}})}\left(32+96\log\frac{T_{\rm{eq}}}{T_{\rm{ew}}}\right)\frac{M_{\rm{Pl}}T_{\rm{R}}^{2}}{\zeta M_{s}^{2}M_{F}}|Q_{\rm G}^{\mathrm{init}}|^{-3/4}&(T_{\rm{R}}<T_{\rm{ew}}<T_{\mathrm{eq}})\vspace{1mm}\\ -\frac{16\pi}{23\sqrt{5}}\frac{1}{\sqrt{g_{*}(T_{\rm{ew}})}}\left(47-15\frac{T_{\mathrm{ew}}^{2}}{T_{\mathrm{R}}^{2}}+96\log\frac{T_{\rm{eq}}}{T_{\mathrm{R}}}\right)\frac{M_{\mathrm{Pl}}T_{\mathrm{R}}^{2}}{\zeta M_{s}^{2}M_{\rm{F}}}|Q_{\rm G}^{\mathrm{init}}|^{-3/4}&(T_{\mathrm{ew}}<T_{\mathrm{R}}<T_{\rm{eq}})\vspace{1mm}\\ -\frac{16\pi}{23\sqrt{5}}\frac{1}{\sqrt{g_{*}(T_{\rm{ew}})}}\left(45-15\frac{T_{\mathrm{ew}}^{2}}{T_{\rm{eq}}^{2}}+2\frac{T_{\mathrm{eq}}}{T_{\mathrm{R}}}\right)\frac{M_{\mathrm{Pl}}T_{\rm{eq}}^{2}}{\zeta M_{s}^{2}M_{\rm{F}}}|Q_{\rm G}^{\mathrm{init}}|^{-3/4}&(T_{\mathrm{ew}}<T_{\rm{eq}}<T_{\mathrm{R}})\end{cases}\ ,$ (3.19)
where $M_{s}$ is the mass of gauginos involved in the evaporation process.
Here, $T_{\rm{eq}}\sim(M_{s}^{2}M_{\rm{F}})^{1/3}|Q_{G}^{\mathrm{init}}|^{-1/12}$ is the temperature at which the evaporation rate of L-balls [51] and the diffusion rate of leptons evaporated from L-balls [52] becomes equal, and $T_{\rm{ew}}$ is the temperature at electroweak phase transition. We assumed $g_{*}(T)=g_{*}(T_{\rm{ew}})$ for $T>T_{\rm{ew}}$. Then, if $T_{\mathrm{R}}<T_{\mathrm{ew}}<T_{\mathrm{eq}}$, the total baryon asymmetry induced from the L-ball evaporation process can be evaluated as
$\eta_{b}$ $\simeq\eta_{L}\left(\frac{\Delta Q_{b}}{Q_{\rm{G}}^{\rm init}}\right)$
$\simeq-5.5\times 10^{-11}\left(\frac{g_{*}}{80}\right)^{-\frac{1}{4}}\left(\frac{\delta}{10^{-3}}\right)\left(\frac{M_{s}}{10^{4}\,{\rm{GeV}}}\right)^{-2}\left(\frac{m_{3/2}}{1\,{\rm{GeV}}}\right)^{\frac{9}{2}}\left(\frac{M_{\rm{F}}}{10^{6}\,{\rm{GeV}}}\right)^{-6}\left(\frac{T_{\rm{R}}}{10\,{\rm{GeV}}}\right)^{2},$ (3.20)
where $\eta_{L}$ is the lepton-to-entropy ratio and we take $\eta_{L}=L_{\nu_{e}}^{\mathrm{init}}$. We require that this resulting baryon asymmetry does not spoil the BBN, that is, $|\eta_{b}|<|\eta_{b}^{\rm{obs}}|\sim 10^{-10}$, which leads to the upper bound of $T_{\rm{R}}$ in terms of $M_{s},m_{3/2}$, and $M_{\rm{F}}$. On the other hand, the L-ball domination ($f_{\mathrm{Q}}\gg 1$) requires $T_{\mathrm{R}}\gtrsim T_{\mathrm{decay}}$. Thus, using Eq. (3.14) and the upper bound on $T_{\mathrm{R}}$, we obtain
$M_{\rm{F}}\gtrsim 7.3\times 10^{5}\,{\rm{GeV}}\left(\frac{g_{*}(T_{\rm{decay}})}{80}\right)^{-3/40}\left(\frac{|\delta|}{10^{-3}}\right)^{1/10}\left(\frac{m_{3/2}}{1\,{\rm{GeV}}}\right)^{19/20}\left(\frac{M_{s}}{10^{4}\,{\rm{GeV}}}\right)^{-1/5}.$ (3.21)
@@
Figure 6: Constraints on $m_{3/2}$ and $M_{\mathrm{F}}$ in the delayed-type L-ball scenario. We set $M_{s}=10\,{\rm{TeV}}$ in the upper panels and $M_{s}=10^{2}\,{\rm{TeV}}$ in the lower panels. We set $\delta=10^{-2}$ and $\delta=10^{-3}$ in the left and right panels, respectively. The green, violet, and brown lines correspond to $|L_{\nu_{e}}^{\mathrm{init}}|=10^{-4}$, $10^{-3}$, and $5\times 10^{-3}$, respectively. The red-shaded region is excluded because baryon asymmetry is overproduced, $|\eta_{b}|>10^{-10}$, due to the evaporation process of L-balls when we require $T_{\rm{R}}>T_{\rm{decay}}$ (see Eq. (3.21)). The cyan region is excluded because $T_{\rm{decay}}$ is lower than $T_{\rm{low}}$ (see Eq. (3.18)).
@@ 4.1 Free-streaming length of sterile neutrinos
where $m_{\rm{eff}}=94\,{\rm{eV}}\times\Omega_{\rm{DM}}h^{2}$ is the total mass of the active neutrinos for their energy density to be the same as dark matter, and $T_{\nu}$ is the temperature of active neutrinos. Using the relation between $m_{x}$ and $\lambda_{\mathrm{FS}}$, we translate the lower bound, $m_{x}>4.09$ keV ($95\%$ CL) [54], into the upper bound of $\lambda_{\rm{FS}}$, $\lambda_{\rm{FS}}<\lambda_{\rm{FS}}^{\rm{upper}}=0.099$ Mpc.
@@
where we used $p=\epsilon_{\rm{phys}}(T_{*})T_{*}a_{*}/a$. Now, we calculate the free-streaming length of the sterile neutrinos in the SF mechanism using the distribution function obtained in Sec. 2. We show the free-streaming length of the sterile neutrinos as a function of $m_{s}$ for a given $L_{\nu_{e}}^{\rm{init}}$ in Fig. 7. By equating $\lambda_{\rm{FS}}(m_{s},L_{\nu_{e}}^{\rm{init}})$ and $\lambda_{\rm{FS}}^{\rm{upper}}=0.099$ Mpc, we obtain the Lyman-$\alpha$ constraint on $m_{s}$ for a given $L_{\nu_{e}}^{\rm{init}}$. We show the lower bounds of $m_{s}$ for several values of $L_{\nu_{e}}^{\rm{init}}$ in Table 1.
@@ 4.2 Gravitational wave enhancement at L-ball decay
A 0.5 $1.1\times 10^{6}$ $1.0\times 10^{2}$ $1.5\times 10^{4}$ $8.7\times 10^{-11}$ $7.0\times 10^{-4}$
B 0.4 $8.5\times 10^{5}$ $1.0\times 10^{3}$ $1.7\times 10^{5}$ $8.7\times 10^{-11}$ $8.4\times 10^{-4}$
C 2.0 $5.4\times 10^{6}$ $1.4\times 10^{4}$ $2.1\times 10^{5}$ $8.7\times 10^{-11}$ $2.2\times 10^{-4}$
Table 2: Benchmark parameters to evaluate the gravitational wave spectra.